OpenStudy (anonymous):

x^4-17x^2+16=0 has how many positive and negative real zeroes?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(x^2-16)(x^2-1)=0 x^2=16,1 x=+-4.+-1

OpenStudy (anonymous):

two _ve real zero and 2 negative

OpenStudy (angela210793):

u can replace x^2 with a t and u'll have t^2-17t+16=0 and u can solve it as a second grade eq.

OpenStudy (anonymous):

Angela knows her stuff.